OpenStudy (jchick):

PLEASE HELP I ALWAYS GET LOST WITH THESE Which equation below is not an equation that represents the graph shown below? line through the points 0 comma 0 and 2 comma 1

OpenStudy (jchick):

a. y - 3 = 2(x - 6) b. y + 3 = one half(x + 6) c. 2y = x d. x - 2y = 0

Nnesha (nnesha):

\[(0 , 0 )( 2,1)\]

Nnesha (nnesha):

use slope formula to find slope or you can replace x and y one by one and see if yo get both side equal then that is your right answer

OpenStudy (anonymous):

First determine the equation of the line, from there, determine what is asked

Nnesha (nnesha):

i guess you should use slop formula to find slope instead of replacing x and y values that would be easy \[(y - y_1 )= m (x-x_1)\] point slope form where m is slope y= mx + b <--- slope intercept form where m is slope which is always with x and b is y-intercept

OpenStudy (jchick):

x- 2y = 0?

Nnesha (nnesha):

\[\huge\rm\frac{ y_2 - y_1 }{ x_2 - x_1 }\] formula to find slope use that two order pair (0,0) and (2,1) where \[\large\rm (x_1 ,y_1) (x_2 , y_2)\]

Nnesha (nnesha):

yep!!!

Nnesha (nnesha):

wait there is another equation with 1/2 slope

Nnesha (nnesha):

how do you know its D ? :)

OpenStudy (jchick):

It was the only one that fit

OpenStudy (jchick):

I graphed them all

Nnesha (nnesha):

yes right :)